# Injective flows for star-like manifolds

### v2

@@ 1 Introduction
Normalizing Flows (NFs) are flexible and efficient models that allow us to accurately estimate arbitrary probability distributions. The key idea is to transform a simple base distribution into a complicated one through a series of bijections. However, in many applications we know that the target density lies on a certain lower-dimensional manifold. A common approach is to have the base distribution defined on the lower-dimensional space and use an injective transformation to embed it into higher dimensions. Unfortunately, the computation of the transformed density involves an expensive Jacobian determinant term, which makes the model computationally prohibitive. Currently, exact and efficient Jacobian determinant is possible only for trivial manifolds like spheres and tori (Gemici et al., 2016; Rezende et al., 2020) or for very restrictive transformations (Ross and Cresswell, 2021). In practice, most work renounce exact density estimation and approximate the Jacobian determinant term (Kumar et al., 2020; Caterini et al., 2021; Sorrenson et al., 2024), often with high variance estimators. Exact density estimation might not be critical when training with maximum likelihood, but is crucial in many applications like variational inference, where samples are not available.
@@ 2 Background
Let $\bm{x}$ be a $d$-dimensional random variable with unknown distribution $p_{x}(\bm{x})$ and let $\bm{z}$ be a $d$-dimensional random variable with known base distribution $p_{z}(\bm{z})$. The key idea of NFs is to model the unknown distribution $p_{x}(\bm{x})$ through a transformation $\operatorname{\mathcal{T}}:\operatorname{\mathbb{R}}^{d}\mapsto\operatorname{\mathbb{R}}^{d}$ such that $\bm{x}=\operatorname{\mathcal{T}}(\bm{z})$. If $\operatorname{\mathcal{T}}$ is a diffeomorphism, i.e. differentiable bijection with differentiable inverse $\operatorname{\mathcal{T}}^{-1}$, the change of variable formula (Rudin, 1987) allows us to express $p_{x}(\bm{x})$ solely in terms of the base distribution $p_{z}(\bm{z})$ and $\operatorname{\mathcal{T}}$: $p_{x}(\bm{x})=p_{z}(\bm{z})\left|\det J_{\operatorname{\mathcal{T}}}(\bm{z})\right|^{-1}$, where $J_{\operatorname{\mathcal{T}}}$ is the Jacobian of the transformation $\operatorname{\mathcal{T}}$. Therefore, the trade-off consists of implementing bijections with tractable $\det J_{\operatorname{\mathcal{T}}}$ which are still flexible enough to approximate any well-behaved distribution. One key idea is to exploit the property that, given a set of bijections $\{\operatorname{\mathcal{T}}^{(i)}\}_{i=1}^{k}$, their composition $\operatorname{\mathcal{T}}=\operatorname{\mathcal{T}}^{(k)}\circ\cdots\circ\operatorname{\mathcal{T}}^{(1)}$ is still a bijection. Since for bijections the Jacobian is a square matrix, the determinant of a composition of bijections factorizes as the product of the determinants of the individual bijections. Overall, NFs are built as
@@
NFs are limited by the use of bijections, which prevents modeling densities on lower dimensional manifolds. In such cases the target distribution lives on a $m$-dimensional manifold $\operatorname{\mathcal{M}}$ embedded in a $d$-dimensional Euclidean space $\operatorname{\mathcal{M}}\subset\operatorname{\mathbb{R}}^{d}$, where $m<d$. In order to constrain $p_{x}(\bm{x})$ to live on the manifold $\operatorname{\mathcal{M}}$, we need an injective transformation that inflates the dimensionality $\operatorname{\operatorname{\mathcal{T}}_{m\rightarrow d}}:\operatorname{\mathbb{R}}^{m}\mapsto\operatorname{\mathbb{R}}^{d}$. The transformed probability distribution $p_{x}(\bm{x})$ can still be computed by accounting for the volume change (Ben-Israel, 1999; Krantz and Parks, 2008):
@@ 3.2 Proposed injective flows for star-like manifolds
We now show how to define injective flows on star-like manifolds such that the Jacobian determinant, and hence the log density, can be evaluated exactly and efficiently. Relevantly, the modeled distribution is expressed in Cartesian coordinates, which is crucial for most applications. In order to do so we compose three transformations $\operatorname{\mathcal{T}}\coloneqq\operatorname{\mathcal{T}_{s\rightarrow c}}\circ\operatorname{\mathcal{T}}_{r}\circ\operatorname{\mathcal{T}}_{\theta}$ (see Figure 2): (i) an arbitrary diffeomorphism that maps to $d-1$ spherical angles $\operatorname{\mathcal{T}}_{\theta}$, (ii) the injective transformation $\operatorname{\mathcal{T}}_{r}$ that parametrizes the radius $r(\bm{\theta})$ as a function of the angles and (iii) the coordinate transformation $\operatorname{\mathcal{T}_{s\rightarrow c}}$ from spherical to Cartesian coordinates. We refer to Appendix A.1 and  A.4 for the explicit expression of the latter and a discussion of its special structure, respectively. This way the resulting densities will be defined by construction on the manifold parametrized by $\operatorname{\mathcal{T}}_{r}$, while the density will be conveniently expressed in Cartesian coordinates. In practice, we can increase the flexibility of $\operatorname{\mathcal{T}}_{\theta}$ by stacking any bijective layers of choice, as shown in Figure 2.
@@
In variational inference settings we assume the manifold to be given. This means that we should have access to the parametrization $r(\bm{\theta})$, which is often trivial (see e.g. Appendix A.4 Eq. (25) for the l-p norm ball and Eq. (26) for the probabilistic simplex). Even if outside the scope of our work, one could train the proposed injective flows on data via maximum likelihood and learn the manifold alongside the density. Learning $r(\bm{\theta})$ remains only possible if the data is on a $d-1$ dimensional manifold, while manifold learning often aims for more dimensional reduction  (Cayton, 2005). By making $r(\bm{\theta})$ learnable, we would also obtain an explicit parametrization of the learnt manifold.
3.3 Limitations
@@ 5 Applications
In this section we showcase some interesting applications of the proposed injective flows. We consider variational inference tasks (see Eq. (27)) where exact density estimation is known to be beneficial. In particular, we overcome some limitations of current work with exact density estimation, which are applicable only to trivial manifolds like spheres or for very restrictive transformations. In Section 5.1 we first illustrate how the proposed approach learns distributions on a simple 3D manifold. In Section 5.2 we use injective flows to define a novel Objective Bayes approach to penalized likelihood problems. Lastly, in Section 5.3 we introduce a general framework for variational inference in probabilistic mixing models, where we constrain the posterior on the probabilistic simplex.
@@ 5.1 Illustrative distributions in 3D
We first illustrate the proposed model on a simple 3D setting. In particular, we train the injective flow in a variational inference setting where the target distribution is known only up to a constant and no samples are given. Note that this is much more challenging than the usual maximum likelihood settings where flows are trained on data. In particular, it is not trivial to learn densities with many modes. To showcase the effectiveness of the proposed method we show the modeled density for increasingly difficult targets: (i) von Mises-Fisher distribution ($\kappa=5$), (ii) mixture of 50 von Mises-Fisher distributions arranged on a spiral ($\kappa=50$) and (iii) sinusoidal checkerboard ($\log\rho(\theta,\phi)=2\sin(6\theta)\sin(6\phi)$), where $\kappa$ is the concentration parameter and $\theta$ and $\phi$ are the polar and azimuthal angles, respectively. In Figure 3 we show that the modeled densities perfectly match the ground truth in all cases. Furthermore, we compare the modeled density with the ground truth over 10’000 samples and obtained an accurate reconstruction: $\textrm{MSE}=0.013$ (a), $0.011$ (b), and $0.024$ (c).
Figure 3: Reconstructed (“ours”) and ground truth (“gt”) density for the proposed injective flow trained via reverse KL divergence without samples. Only the (unnormalized) target is given: (a) von Mises-Fisher ($\kappa=5$), (b) mixture of 50 von Mises-Fisher ($\kappa=50$) and (c) sinusoidal checkerboard.
@@ 5.3 Variational Inference for probabilistic mixing models
Figure 6: Trained injective flow on the simplex with target $p(\bm{\pi})|\mathcal{D})\propto p(\mathcal{D}|\bm{\pi})$ (“uniform” prior, implicitly) and $p(\bm{\pi})|\mathcal{D})\propto p(\mathcal{D}|\bm{\pi})\textrm{Dir}(\bm{\pi})$ (“Dirichlet” prior). Above: we compare the distribution of non-zero entries in posterior samples for 4 fixed likelihood values (so at fixed reconstruction). Below: we compare sparsity patterns of samples $\bm{\pi}$ with uniform (blue) and Dirichlet (orange) prior.
@@ 6 Conclusions
Previous work on injective flows on manifolds relies on approximations or lower bounds to circumvent the computation of the Jacobian determinant term. In this work we showed how to exactly and efficiently compute the Jacobian determinant term for the general class of star-like manifolds. We then showed that the proposed flow allows for interesting applications that were not possible before. First, with the proposed framework we introduced a novel Objective Bayes approach to penalized likelihood methods. The idea is to circumvent the choice of a subjective prior by constraining the posterior on the manifold defined by level-sets of the prior. Second, we introduced a general variational inference framework for modeling the posterior in probabilistic mixing models. Overall, the proposed framework allows us to efficiently model distributions on arbitrary star-like manifolds and to flexibly specify any choice of prior and likelihood.
@@ A.5 Implementation details
  • bijective layers $\operatorname{\mathcal{T}}_{z}$ and $\operatorname{\mathcal{T}}_{\theta}$. The first bijection $\operatorname{\mathcal{T}}_{z}:\bm{z}\mapsto\bm{z^{\prime}}$ consists of arbitrary (conditional) bijective layers conditioned on the parameter $\lambda$. The conditioning is realized with an expressive Residual network. Then, $\operatorname{\mathcal{T}}_{\theta}:\bm{z^{\prime}}\mapsto\bm{\theta}$ maps the transformed $\bm{z^{\prime}}$ into spherical angles $\bm{\theta}\in U_{\theta}^{d-1}$. This last transformation is also a bijection that can be implemented with an element-wise non-linear activation like Sigmoid (hence diagonal Jacobian). Otherwise, one could use a base distribution which is already defined on the $d-1$ spherical angles and use a bijective transformation that transforms $\bm{\theta}$ within their domain $U_{\theta}^{d-1}$ as $\operatorname{\mathcal{T}}_{\textrm{circ}}:\bm{\theta}\in U_{\theta}^{d-1}\mapsto\bm{\theta^{\prime}}\in U_{\theta}^{d-1}$. We use the circular bijective layers proposed by Rezende et al. [2020] because they allow to nicely integrate the boundary conditions arising from the use of spherical coordinates. In particular, circular layers automatically enforce continuity of the density at the boundary of the domain. Circular layers require the base distribution to be defined on $U_{\theta}^{d-1}$. In practice, we use the distribution of spherical angles, which results in uniform points on the $d-1$ dimensional sphere, and can be implemented efficiently. We use the implementation of circular layers provided in Stimper et al. [2023].

### v3

@@ 1 Introduction
Normalizing Flows (NFs) are flexible and efficient models that allow us to accurately estimate arbitrary probability distributions. The key idea is to transform a simple base distribution into a complicated one through a series of bijections. However, in many applications we know that the target density lies on a certain lower-dimensional manifold. A common approach is to have the base distribution defined on the lower-dimensional space and use an injective transformation to embed it into higher dimensions. Unfortunately, the computation of the transformed density involves an expensive Jacobian determinant term, which makes the model computationally prohibitive. Currently, exact and efficient Jacobian determinant is possible only for trivial manifolds like spheres and tori (Gemici et al., 2016; Rezende et al., 2020) or for very restrictive transformations (Ross and Cresswell, 2021). In practice, most work renounce exact density estimation and approximate the Jacobian determinant term (Kumar et al., 2020; Kothari et al., 2021; Sorrenson et al., 2024), often with high variance estimators. Exact density estimation might not be critical when training with maximum likelihood, but is crucial in many applications like variational inference, where samples are not available.
@@ 2 Background
Let $\bm{x}$ be a $d$-dimensional random variable with unknown distribution $p_{x}(\bm{x})$ and let $\bm{z}$ be a $d$-dimensional random variable with known base distribution $p_{z}(\bm{z})$. The key idea of NFs is to model the unknown distribution $p_{x}(\bm{x})$ through a transformation $\operatorname{\mathcal{T}}:\operatorname{\mathbb{R}}^{d}\mapsto\operatorname{\mathbb{R}}^{d}$ such that $\bm{x}=\operatorname{\mathcal{T}}(\bm{z})$. If $\operatorname{\mathcal{T}}$ is a diffeomorphism, i.e. differentiable bijection with differentiable inverse $\operatorname{\mathcal{T}}^{-1}$, the change of variable formula (Rudin, 1987) allows us to express $p_{x}(\bm{x})$ solely in terms of the base distribution $p_{z}(\bm{z})$ and $\operatorname{\mathcal{T}}$: $p_{x}(\bm{x})=p_{z}(\bm{z})\left|\det J_{\operatorname{\mathcal{T}}}(\bm{z})\right|^{-1}$, where $J_{\operatorname{\mathcal{T}}}$ is the Jacobian of the transformation $\operatorname{\mathcal{T}}$. The hypothesis can be relaxed to $\operatorname{\mathcal{T}}$ being a diffeomorphism a. e. (Munkres and Spivak, 1965), i.e. except on zero-measure sets. Therefore, the trade-off consists of implementing bijections with tractable $\det J_{\operatorname{\mathcal{T}}}$ which are still flexible enough to approximate any well-behaved distribution. One key idea is to exploit the property that, given a set of bijections $\{\operatorname{\mathcal{T}}^{(i)}\}_{i=1}^{k}$, their composition $\operatorname{\mathcal{T}}=\operatorname{\mathcal{T}}^{(k)}\circ\cdots\circ\operatorname{\mathcal{T}}^{(1)}$ is still a bijection. Since for bijections the Jacobian is a square matrix, the determinant of a composition of bijections factorizes as the product of the determinants of the individual bijections. Overall, NFs are built as
@@
NFs are limited by the use of bijections, which prevents modeling densities on lower dimensional manifolds. In such cases the target distribution lives on a $m$-dimensional manifold $\operatorname{\mathcal{M}}$ embedded in a $d$-dimensional Euclidean space $\operatorname{\mathcal{M}}\subset\operatorname{\mathbb{R}}^{d}$, where $m<d$. In order to constrain $p_{x}(\bm{x})$ to live on the manifold $\operatorname{\mathcal{M}}$, we need an injective transformation that inflates the dimensionality $\operatorname{\operatorname{\mathcal{T}}_{m\rightarrow d}}:\operatorname{\mathbb{R}}^{m}\mapsto\operatorname{\mathbb{R}}^{d}$. The transformed probability distribution $p_{x}(\bm{x})$ can still be computed by accounting for the volume change (Ben-Israel, 1999):
@@ 3.2 Proposed injective flows for star-like manifolds
We now show how to define injective flows on star-like manifolds such that the Jacobian determinant, and hence the log density, can be evaluated exactly and efficiently. Relevantly, the modeled distribution is expressed in Cartesian coordinates, which is crucial for most applications. In order to do so we compose three transformations $\operatorname{\mathcal{T}}\coloneqq\operatorname{\mathcal{T}_{s\rightarrow c}}\circ\operatorname{\mathcal{T}}_{r}\circ\operatorname{\mathcal{T}}_{\theta}$ (see Figure 2): (i) an arbitrary diffeomorphism that maps to $d-1$ spherical angles $\operatorname{\mathcal{T}}_{\theta}$, (ii) the injective transformation $\operatorname{\mathcal{T}}_{r}$ that parametrizes the radius $r(\bm{\theta})$ as a function of the angles and (iii) the coordinate transformation $\operatorname{\mathcal{T}_{s\rightarrow c}}$ from spherical to Cartesian coordinates. Note that the change of coordinates $\operatorname{\mathcal{T}_{s\rightarrow c}}$ is a diffeomorphism almost everywhere, i.e. except on a zero-measure set. Therefore, the change of variable formula holds and the probability resulting from the transformation is still exact. We refer to Appendix A.1 and  A.4 for the explicit expression of $\operatorname{\mathcal{T}_{s\rightarrow c}}$ and a discussion of its special structure, respectively. This way the resulting densities will be defined by construction on the manifold parametrized by $\operatorname{\mathcal{T}}_{r}$, while the density will be conveniently expressed in Cartesian coordinates. In other words, the push-forward samples are points in the ambient space.
@@
3.3 Further details and limitations
Variational inference vs maximum likelihood
In this paper we focus on variational inference settings (without observations), and not on the maximum likelihood setting (with observations), which is the main focus of most injective flows. Our motivation is two-fold. First, our method is designed for star-like manifolds (with codimension 1), which are more useful in variational inference settings. In maximum likelihood applications, the underlying manifolds are often assumed to much lower dimensional. Second, while in maximum likelihood settings the approximate Jacobian work well in practice, in variational inference settings the exact Jacobian determinant is crucial to learn the correct target distribution. We show this empirically, already in very simple cases.
Implementation details
The proposed approach is straightforward to implement and only requires to adapt the function $r(\bm{\theta})$ to the star-like manifold under study. This means that we should have access to the parametrization $r(\bm{\theta})$, which is often trivial (see e.g. Appendix A.4 Eq. (25) and Eq. (26)). Our method is also versatile as the transformation $\mathcal{T}_{z}$ can be implemented with any bijective layers of choice. Lastly, we easily avoid numerical instabilities arising from singular points in $\operatorname{\mathcal{T}_{s\rightarrow c}}$ by offsetting the critical angles with a small epsilon.
Limitations
@@ 5 Applications
In this section we showcase some interesting applications of the proposed injective flows. In particular, we overcome some limitations of current work with exact density estimation, which are applicable only to trivial manifolds or for very restrictive transformations. In Section 5.1 we show empirically that the proposed approach provides a significant speedup compared to the explicit computation of the Jacobian determinant. We also show that in variational inference the exact Jacobian determinant is crucial for training while the approximation commonly employed in injective flows results in poor reconstruction. In Section 5.2 we illustrate how the proposed approach learns distributions on simple 3D manifolds. In Section 5.3 we use injective flows to define a novel Objective Bayes approach to penalized likelihood problems. Lastly, in Section 5.4 we introduce a general framework for variational inference in probabilistic mixing models.
5.1 Effectiveness and efficiency of the proposed method
We showed that for star-like manifolds we can efficiently compute the Jacobian determinant in Eq. (2) and we argued that exact evaluation is crucial for variational inference. We now verify both statements empirically. Firstly, we compare the runtime associated with computing the Jacobian determinant via the explicit formula in Eq. (2) with our approach in Eq. (4). In particular, we use a very simple manifold, the hypersphere in $d$ dimensions, and measure runtime over 20 repetitions of Jacobian determinant computation. The results in Figure 3a show that our method provides a significant speedup compared to the explicit computation of the Jacobian. By fitting a linear function to the log-runtime and log-dimension we get that the explicit computation is approximately cubic, $O(d^{2.96})$, while our proposed approach is approximately quadratic, $O(d^{1.81})$. Secondly, we show that training with the exact Jacobian determinant is crucial for variational inference. To do so we first train the proposed injective flow with the exact (and efficient) Jacobian in Eq. (4). Then, we train a second identical flow where the gradient of the Jacobian is approximated with the Hutchinson trace estimator with different number of Gaussian samples $n=1,10,50,100$. This approach is common in most injective flow papers (Caterini et al., 2021; Flouris and Konukoglu, 2023). We use the implementation of state-of-the-art work (Sorrenson et al., 2024). The task is to learn the uniform distribution on a lp-(pseudo) norm ball with $p=0.5$. In Figure  3 we compare the ground truth (log-) density with that obtained with the two models. Note that at test time the model trained with the approximate Jacobian is evaluated with the exact Jacobian. In the Appendix in Figure 10 we measure the similarity between samples of the two models and samples from the true distribution. In all cases the proposed model achieves significantly better results both in terms of density reconstruction and samples quality. Even at the cost of increased runtime (shown in Figure 10), increasing the number of samples the approximate method fails to achieve similar performance to ours.
Figure 3: (a) Runtime comparison of Jacobian determinant computation with proposed approach in Eq (4) and with the brute-force computation Eq. (2). (b) MSE of learned log-density for an injective flow trained to learn a uniform distribution on a $l_{p}$-(pseudo) norm ball with $p=0.5$. We compare the same model trained with our exact Jacobian and with the Hutchinson trace estimator with $n=1,10,50,100$ number of Gaussian samples. Note that the number of samples is upper bounded by the dimensionality of the problem and that at evaluation the exact Jacobian is used.
@@ 5.2 Illustrative distributions in 3D
We first illustrate the proposed model on a simple 3D setting. In particular, we train the injective flow in a variational inference setting where the target distribution is known only up to a constant and no samples are given. Note that this is much more challenging than the usual maximum likelihood settings where flows are trained on data. In particular, it is not trivial to learn densities with many modes. To showcase the effectiveness of the proposed method we show the modeled density for increasingly difficult targets: (i) von Mises-Fisher distribution ($\kappa=5$), (ii) mixture of 50 von Mises-Fisher distributions arranged on a spiral ($\kappa=50$) and (iii) sinusoidal density ($\log\rho(\theta,\phi)=\sin(4\theta)\sin(4\phi)$), where $\kappa$ is the concentration parameter and $\theta$ and $\phi$ are the polar and azimuthal angles, respectively. The sinusoidal density (iii) is defined on a deformed sphere. In Figure 4 we show that the modeled densities perfectly match the ground truth in all cases. Furthermore, we compare the modeled density with the ground truth over 10’000 samples and obtained an accurate reconstruction: $\textrm{MSE}=0.013$ (a), $0.011$ (b). As the normalized density on the deformed sphere is not given, we only provide a qualitative visualization, which shows an accurate reconstruction.
Figure 4: Learned (“ours”) and ground truth (“gt”) density for proposed injective flow trained via reverse KL divergence (no samples). Only the (unnormalized) target is given: (a) von Mises-Fisher ($\kappa=5$), (b) mixture of 50 von Mises-Fisher ($\kappa=50$) and (c) sinusoidal density on deformed sphere
@@ 5.4 Variational Inference for probabilistic mixing models
Figure 7: Trained injective flow on the simplex with target $p(\bm{\pi}|\mathcal{D})\propto p(\mathcal{D}|\bm{\pi})$ (“uniform” prior, implicitly) and $p(\bm{\pi}|\mathcal{D})\propto p(\mathcal{D}|\bm{\pi})\textrm{Dir}(\bm{\pi})$ (“Dirichlet” prior). Above: we compare the distribution of non-zero entries in posterior samples for 4 fixed likelihood values (so at fixed reconstruction). Below: we compare sparsity patterns of samples $\bm{\pi}$ with uniform (blue) and Dirichlet (orange) prior.
@@ 6 Conclusions
Previous work on injective flows on manifolds relies on approximations or lower bounds to circumvent the computation of the Jacobian determinant term. In this work we showed how to exactly and efficiently compute the Jacobian determinant term for the general class of star-like manifolds. We validate empirically the claimed computational advantage and we show that exact Jacobian computation is crucial for variational inference, already in very simple settings. We then showed that the proposed flow allows for interesting applications that were not possible before. First, with the proposed framework we introduced a novel Objective Bayes approach to penalized likelihood methods. The idea is to circumvent the choice of a subjective prior by constraining the posterior on the manifold defined by level-sets of the prior. Second, we introduced a general variational inference framework for modeling the posterior in probabilistic mixing models. Overall, the proposed framework allows us to efficiently model distributions on arbitrary star-like manifolds and to flexibly specify any choice of prior and likelihood.
@@ A.5 Implementation details
  • bijective layers $\operatorname{\mathcal{T}}_{z}$ and $\operatorname{\mathcal{T}}_{\theta}$. The first bijection $\operatorname{\mathcal{T}}_{z}:\bm{z}\mapsto\bm{z^{\prime}}$ consists of arbitrary (conditional) bijective layers conditioned on the parameter $\lambda$. The conditioning is realized with an expressive Residual network. Then, $\operatorname{\mathcal{T}}_{\theta}:\bm{z^{\prime}}\mapsto\bm{\theta}$ maps the transformed $\bm{z^{\prime}}$ into spherical angles $\bm{\theta}\in U_{\theta}^{d-1}$. This last transformation is also a bijection that can be implemented with an element-wise non-linear activation like Sigmoid (hence diagonal Jacobian). Otherwise, one could use a base distribution which is already defined on the $d-1$ spherical angles and use a bijective transformation that transforms $\bm{\theta}$ within their domain $U_{\theta}^{d-1}$ as $\operatorname{\mathcal{T}}_{\textrm{circ}}:\bm{\theta}\in U_{\theta}^{d-1}\mapsto\bm{\theta^{\prime}}\in U_{\theta}^{d-1}$. In short, any bijective layers followed by a element-wise non linear function that maps to spherical angles could be used. We use the circular spline layers proposed by Rezende et al. [2020] because they allow to nicely integrate the boundary conditions arising from the use of spherical coordinates. These layers are based on the neural spline layers proposed in Durkan et al. [2019], which consist of a combination of $K$ segments where each segment is a simple rational-quadratic function. The flexibility of the layers increase with $K$. This way the transformation can be designed such that it is monotonically increasing (hence invertible) and such that it fulfills given boundary conditions. The main difference with Durkan et al. [2019] is that circular splines require periodic boundary conditions in order to enforce continuity of the density at the boundary. This way we can define bijections from $[0,2\pi]$ to itself. As a consequence, circular splines require the base distribution to be defined on $U_{\theta}^{d-1}$. In practice, we use the distribution of spherical angles, which results in uniform points on the $d-1$ dimensional sphere, and can be implemented efficiently. We use the implementation of circular layers provided in Stimper et al. [2023]. Such construction scales well with the dimensions.